OpenStudy (anonymous):

HELP!! If the point P is on the graph of a function f, find the corresponding point on the graph of the given function. P(5, 8); y = 1/4f (1/3x) − 1

OpenStudy (anonymous):

we can do this i think

OpenStudy (anonymous):

if i am reading it correctly, it is \(y=\frac{1}{4}f(\frac{1}{3}x)-1\) is that right?

OpenStudy (anonymous):

?? i don't want to begin if that is the wrong thing

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok good' so we only know the value of \(f(5)\) therefore we need to figure out what we would put in for \(x\) in \(f(\frac{1}{3}x)\) to get 5 set \[\frac{1}{3}x=5\] get \(x=15\)

OpenStudy (anonymous):

now if \(x=15\) then \(f(\frac{1}{3}x)=f(\frac{1}{3}\times 15)=f(5)=8\) and so \[\frac{1}{4}\times 8-2=2-1=1\] therefore the "corresponding point" is \((15,1)\)

OpenStudy (anonymous):

thanks i get how to do it now :)